I need to solve for L
p=2L+2w

Answer by ankor@dixie-net.com(22740)

I need to solve for L
p=2L+2w
Or
2L + 2W = P
:
Subtract 2W from both sides and you have:
2L = P - 2W
:
Divide both sides by 2
L = (P-2W)/2
or
L = P/2 - W